OpenStudy (anonymous):

64, 16, 4, 1, whats the common ratio?

OpenStudy (asnaseer):

use the same method described to you by @Outkast3r09 in your earlier question: http://openstudy.com/updates/4fe52d86e4b06e92b87389c1

jimthompson5910 (jim_thompson5910):

If you're still stuck, take any term (but the first one) and divide it by the previous term. This will give you the common ratio.

OpenStudy (anonymous):

so 4?

jimthompson5910 (jim_thompson5910):

you're dividing the wrong way

OpenStudy (anonymous):

ohhh so it's 1/4? but is it negative or positive?

jimthompson5910 (jim_thompson5910):

it's positive since 4/16 = 1/4 or 16/64 = 1/4

jimthompson5910 (jim_thompson5910):

you can also see this because the signs don't change from term to term

OpenStudy (anonymous):

okay can you help me with this one? What is the next term of the following sequence? 12, 6, 3, ... is it going to be 0?

jimthompson5910 (jim_thompson5910):

What are you doing to 12 to get to 6?

OpenStudy (anonymous):

subtracting 6

jimthompson5910 (jim_thompson5910):

what about jumping from 6 to 3?

OpenStudy (anonymous):

ohhh okayyy its going to be 1 and 1/2 right?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

or 3/2

OpenStudy (anonymous):

okayy one more question, What is the next term of the following sequence? 12, 6, 3, ...

jimthompson5910 (jim_thompson5910):

you just asked that

OpenStudy (anonymous):

is it 405?